# On eigenvalues of the Landau Hamiltonian with a periodic electric potential

### v1

@@ 4 The proof of Theorem 7
The functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=2,3$, satisfy properties (1), (2), and (3) as well as functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=0,1$, (for the functions ${\mathcal{F}}_{j}^{(2)}$, property (2) is fulfilled by Lemma 12).
From Lemma 10, it follows that equations (4.10) hold for $\varepsilon\in{\mathbb{C}}$ iff
$\tau=\tau^{(0)}+\frac{\varepsilon}{2}\,\sum\limits_{\mu}\tau_{\mu}(w_{*},{\mathcal{F}}^{(3)}_{\mu})_{L^{2}_{\Lambda(T)}},$ (4.11)
$b_{\nu}=b_{\nu}^{(0)}+\frac{\varepsilon}{2}\,\sum\limits_{\mu}R_{\nu\mu}(w_{*},{\mathcal{F}}^{(3)}_{\mu})_{L^{2}_{\Lambda(T)}},\quad\nu=1,\dots,m,$ (4.12)
@@
$\tau^{(0)}\doteq\frac{1}{2}\,\sum\limits_{\mu}\tau_{\mu}(w_{*},\widetilde{\mathbb{W}}_{\mu})_{L^{2}_{\Lambda(T)}},\quad b_{\nu}^{(0)}\doteq\frac{1}{2}\,\sum\limits_{\mu}R_{\nu\mu}(w_{*},\widetilde{\mathbb{W}}_{\mu})_{L^{2}_{\Lambda(T)}}.$
@@
$\tau^{(j+1)}=\tau^{(0)}+\frac{\varepsilon}{2}\,\sum\limits_{\mu}\tau_{\mu}\big(w_{*},{\mathcal{F}}^{(3)}_{\mu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})\bigr)_{L^{2}_{\Lambda(T)}},$
$b_{\nu}^{(j+1)}=b_{\nu}^{(0)}+\frac{\varepsilon}{2}\,\sum\limits_{\mu}R_{\nu\mu}\big(w_{*},{\mathcal{F}}^{(3)}_{\mu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})\bigr)_{L^{2}_{\Lambda(T)}},$

### v2

@@ 4 The proof of Theorem 7
The functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=2,3$, as well as the functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=0,1$, satisfy properties (1), (2), and (3) (for the functions ${\mathcal{F}}_{j}^{(2)}$, property (2) is fulfilled by Lemma 12).
Since $\|w_{*}\|^{2}_{L^{2}_{\Lambda(T)}}=2T$, it follows from Lemma 10 that equations (4.10) hold for $\varepsilon\in{\mathbb{C}}$ iff
$\tau=\tau^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}\tau_{\mu}(w_{*},{\mathcal{F}}^{(3)}_{\mu})_{L^{2}_{\Lambda(T)}},$ (4.11)
$b_{\nu}=b_{\nu}^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}R_{\nu\mu}(w_{*},{\mathcal{F}}^{(3)}_{\mu})_{L^{2}_{\Lambda(T)}},\quad\nu=1,\dots,m,$ (4.12)
@@
$\tau^{(0)}\doteq\frac{1}{2T}\,\sum\limits_{\mu}\tau_{\mu}(w_{*},\widetilde{\mathbb{W}}_{\mu})_{L^{2}_{\Lambda(T)}},\quad b_{\nu}^{(0)}\doteq\frac{1}{2T}\,\sum\limits_{\mu}R_{\nu\mu}(w_{*},\widetilde{\mathbb{W}}_{\mu})_{L^{2}_{\Lambda(T)}}.$
@@
$\tau^{(j+1)}=\tau^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}\tau_{\mu}\big(w_{*},{\mathcal{F}}^{(3)}_{\mu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})\bigr)_{L^{2}_{\Lambda(T)}},$
$b_{\nu}^{(j+1)}=b_{\nu}^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}R_{\nu\mu}\big(w_{*},{\mathcal{F}}^{(3)}_{\mu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})\bigr)_{L^{2}_{\Lambda(T)}},$